[troffman,twoside]article document pbm521 September 1989

NAME pbm - portable bitmap file format

DESCRIPTION The portable bitmap format is a lowest common denominator monochrome file format. It was originally designed to make it reasonable to mail bitmaps between different types of machines using the typical stupid network mailers we have today. Now it serves as the common language of a large family of bitmap conversion filters. The definition is as follows: IPlist - A “magic number” for identifying the file type. A pbm file’s magic number is the two characters “P1”. - Whitespace (blanks, TABs, CRs, LFs). - A width, formatted as ASCII characters in decimal. - Whitespace. - A height, again in ASCII decimal.

- Whitespace. - Width * height bits, each either ’1’ or ’0’, starting at the top-left corner of the bitmap, proceding in normal English reading order. - The character ’1’ means black, ’0’ means white. - Whitespace in the bits section is ignored. - Characters from a “#” to the next end-of-line are ignored (comments). - No line should be longer than 70 characters.